OpenStudy (bookworm14):

Can you check my answers please? These sort of confused me, but I might have it right. Thanks! (will post below just gimme a sec :D )

OpenStudy (bookworm14):

Okay, I wrote my answers below each question, please lemme know If you cannot see or understand my answers ni the pic, thanks!

OpenStudy (lovedance11):

Nope

OpenStudy (smilie2015):

2lw+2lh+2wh=s w(2l+2h)+2lh=s w(2l+2h)+2lh-s=0 2lh-s=w(2l+2h) 2lh-s/2l+2h=w

OpenStudy (johnweldon1993):

^Not quite, if you wanted to do it that way you forgot a negative sign

OpenStudy (smilie2015):

ur second answer is right

OpenStudy (johnweldon1993):

\[\large s = 2(lw + lh + wh)\] \[\large s = 2lw + 2lh + 2wh\] \[\large s - 2lh = 2lw + 2wh\] \[\large s - 2lh = w(2l + 2h)\] \[\large w = \frac{s - 2lh}{2l + 2h}\]

OpenStudy (bookworm14):

ohhh I forgot to put my plus sign at the bottom when I wrote it, haha my bad, thank you :)

OpenStudy (johnweldon1993):

Yeah exactly, it is so small and obvious that I overlooked it as well in the beginning, but glad you caught it ^_^

OpenStudy (bookworm14):

Oh yeah, boom! :D I did it right , so freaking happy right now :DDD Thank you John!

OpenStudy (johnweldon1993):

Of course love :D